Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 586
Entry time: 12 Dec 2018, 10:17
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 09.30 |
|
|
SPB/SFX
- Radiation damage tests were successful.
- Instrument is set for the power cut.
FXE
- According to the information available to AE the instrument should be set up.
SQS
- Successful users run in the last 2 weeks
- Some interlocks need to be activated. Discussion with AE ongoing
- Loops updates to be discussed/executed plus CR status will be checked
- Triggers requests are being sicussed with BF
SCS
MID
HED
- Loop 1 will be operative probably today. CAS support for creation of scenes has been asked.
- Cabling ongoing.
- Additional Racks in A27 have been installed
OPT Lasers
- As soon as the system is ready for the power cut in SASE1 a mail will be sent to PSPO.
XRO
XPD
VAC
AE
- XTD6 crate 219 has to be rebooted.
EET
ITDM
- Maintenance and updates (e.g. online clusters ) has started as indicated in the shutdown activities planning. Should go on until the end of the week.
DET
CAS
- 2.3.1. will be deployed in SASE3 tunnels today
- Schedule for 2019 has been circulated by HS to the readiness mailing list
AoB
|